Nature of Uncollectible Accounts

The XYZ Corporation owns and operates hotels and casinos including the XYZ Grand and the Bellagio in Las Vegas, Nevada. As of a recent year, XYZ reported accounts receivable of $464,000 and allowance for doubtful accounts of $58,464.

Patient Care manufactures and sells a wide range of healthcare products including Band-Aids and Tylenol. As of a recent year, Patient Care reported accounts receivable of $886,000 and allowance for doubtful accounts of $29,238.

Round your answers to one decimal place

a. Compute the percentage of the allowance for doubtful accounts to the accounts receivable for The XYZ Corporation.

b. Compute the percentage of the allowance for doubtful accounts to the accounts receivable for Patient Care.
Answer #1

a) Percentage of the allowance for doubtful account for XYZ corporation = 58464/464000 = 12.60%

b) Percentage of the allowance for doubtful account for Patient Care = 29238/886000 = 3.3%